OpenStudy (sophia67):

Given f(x) =the quantity of 5x plus 1, divided by 2 , solve for f−1(8)

OpenStudy (eliesaab):

Is this your function \[ f(x)=\frac{5x+1}2 \]

OpenStudy (eliesaab):

and you want \[ f^{-1}(8) \]

OpenStudy (eliesaab):

Solve for x \[ \frac{5x+1}2=8 \] and you will be done

OpenStudy (eliesaab):

Or 5 x+1 =16 What is x?